# Matryoshka: Learning to Drive Black-Box LLMs with LLMs

### v1

@@ G.2 GSM-Hard
For Matryoshka, we employ a three-shot prompt to guide the decomposition steps, where <question> represents the current problem.
 <problem_decomposition> Prompt
 # System Message: You will decompose a math problem into smaller
 parts. Follow the prompt instruction and do not generate redundant
 information.
 Q: Olivia has \$23. She bought five bagels for \$3 each. How much money
 does she have left?
@@
 Q: Michael had 58 golf balls. On tuesday, he lost 23 golf balls. On
 wednesday, he lost 2 more. How many golf balls did he have at the end
 of wednesday?
@@
[…]
 Michael have left after losing the total amount?
 Q: There were nine computers in the server room. Five more computers
 were installed each day, from monday to thursday. How many computers
 are now in the server room?
[…]
 server room after adding the new ones?
 Q: <question>
 A:
Code Generation.
Given the problem and the decomposition provided by Matryoshka, the Black-box model generates the corresponding code block for each sub-problem. We continue to use a three-shot prompt to instruct the Black-box model on how to translate the sub-problems into code, where <question> represents the current problem and <decompose> represents the decomposition provided by Matryoshka.
 <code_generation> Prompt
 # System Message: You will write python program to solve math
 problems. You will write annotations and code blocks following
 instructions. Annotations should be written in the form of a question.
 Let’s use python to solve math problems. Here are three examples how
 to do it,
 Q: Olivia has \$23. She bought five bagels for \$3 each. How much money
 does she have left?
[…]
 def solution():
     """Olivia has \$23. She bought five bagels for
     \$3 each. How much money does she have left?"""
     # Initialization of variables
     money_initial = 23
     bagels = 5
     bagel_cost = 3
     # How much does Olivia spend on bagels?
     money_spent = bagels * bagel_cost
     # How much money does Olivia have left after the purchase?
     money_left = money_initial - money_spent
     result = money_left
     return result
 Q: Michael had 58 golf balls. On tuesday, he lost 23 golf balls. On
 wednesday, he lost 2 more. How many golf balls did he have at the end
 of wednesday?
[…]
 have left after losing the total amount?
 def solution():
     """Michael had 58 golf balls. On tuesday, he lost 23 golf balls.
     On wednesday, he lost 2 more. How many golf balls did he have at
     the end of wednesday?"""
     # Initialization of variables
     golf_balls_initial = 58
     golf_balls_lost_tuesday = 23
     golf_balls_lost_wednesday = 2
     # How many golf balls did Michael lose in total by the end of
     Wednesday?
     golf_balls_left = golf_balls_initial - golf_balls_lost_tuesday -
     golf_balls_lost_wednesday
     # How many golf balls does Michael have left after losing the
     total amount?
     result = golf_balls_left
     return result
 Q: There were nine computers in the server room. Five more computers
 were installed each day, from monday to thursday. How many computers
 are now in the server room?
[…]
 room after adding the new ones?
 def solution():
     """There were nine computers in the server room. Five more
     computers were installed each day, from monday to thursday. How
     many computers are now in the server room?"""
     # Initialization of variables
     computers_initial = 9
     computers_per_day = 5
     num_days = 4  # 4 days between monday and thursday
     # How many computers were added in total from Monday to Thursday?
     computers_added = computers_per_day * num_days
     # How many computers are now in the server room after adding the
     new ones?
     computers_total = computers_initial + computers_added
     result = computers_total
     return result
 How about this question?
 Q: <question>
 <decompose>
@@
 <refinement> Prompt
 Let’s use python to solve math problems. Here are three successful
 cases on how to do it,
 Q: Olivia has \$23. She bought five bagels for \$3 each. How much money
 does she have left?
 def solution():
     """Olivia has \$23. She bought five bagels for \$3 each. How much
     money does she have left?"""
     money_initial = 23
     bagels = 5
     bagel_cost = 3
     money_spent = bagels * bagel_cost
     money_left = money_initial - money_spent
     result = money_left
     return result
 Q: Michael had 58 golf balls. On tuesday, he lost 23 golf balls. On
 wednesday, he lost 2 more. How many golf balls did he have at the end
 of wednesday?
 def solution():
     """Michael had 58 golf balls. On tuesday, he lost 23 golf balls.
     On wednesday, he lost 2 more. How many golf balls did he have at
     the end of wednesday?"""
     golf_balls_initial = 58
     golf_balls_lost_tuesday = 23
     golf_balls_lost_wednesday = 2
     golf_balls_left = golf_balls_initial - golf_balls_lost_tuesday -
     golf_balls_lost_wednesday
     result = golf_balls_left
     return result
 Q: There were nine computers in the server room. Five more computers
 were installed each day, from monday to thursday. How many computers
 are now in the server room?
 def solution():
     """There were nine computers in the server room. Five more
     computers were installed each day, from monday to thursday. How
     many computers are now in the server room?"""
     computers_initial = 9
     computers_per_day = 5
     num_days = 4  # 4 days between monday and thursday
     computers_added = computers_per_day * num_days
     computers_total = computers_initial + computers_added
     result = computers_total
     return result
 # Here is the actual question.
 Q: <question>
 You have generated code of solution() to solve the task. However, you
 executed the solution() function and get an error message:
 <error_msg>
 Referring to the successful case and the error message, you should
 complete the solution function with the correct code.
G.3 LaMP
Following the RAG-based framework (Salemi et al., 2023) and the PAG-based framework (Richardson et al., 2023), we implement prompt designs for both Matryoshka and the baseline methods. The prompt design for RAG is presented in Table 7, while the prompts for the two-stage PAG and Matryoshka are shown in Table 8. We create prompts for the controller model using the templates from Table 8 and subsequently combine the intermediate generations with the input question to form prompts for the environment model. Since LaMP-3 prompts are particularly lengthy, we provide additional examples of our PAG prompts for LaMP-1, LaMP-2N, LaMP-2M, and LaMP-4 as follows.
@@ G.3 LaMP
Task Per Profile Entry Prompt (PPEP) Aggregated Input Prompt (AIP) LaMP-1 "Pi[title]" concat([PPEP(P1), …, PPEP(Pn)], ", and "). [INPUT] LaMP-2N "the category for the article: "Pi[text]" is ""Pi[category]"" concat([PPEP(P1), …, PPEP(Pn)], ", and "). [INPUT] LaMP-2M "the tag for the movie: "Pi[description]" is "Pi[tag]" concat([PPEP(P1), …, PPEP(Pn)], ", and "). [INPUT] LaMP-3 Pi[score] is the score for "Pi[text]" concat([PPEP(P1), …, PPEP(Pn)], ", and "). [INPUT] LaMP-4 "Pi[title]" is the title for "Pi[text]" concat([PPEP(P1), …, PPEP(Pn)], ", and "). [INPUT]
@@
Task Prompt LaMP-1 Write a summary, in English, of the research interests and topics of a researcher who has published the following papers. Only generate the summary, no other text. LaMP-2N Look at the following past articles this journalist has written and determine the most popular category they write in. Answer in the following format: most popular category: <category top1>, <category top2>, …, <category topn> LaMP-2M Look at the following past movies this user has watched and determine the mostpopular tag they labeled. Answer in the following form: most popular tag: <tag top1>, <tag top2>, …, <tag topn> LaMP-3 Based on this user’s past reviews, what are the most common scores they give for positive and negative reviews? Answer in the following form: most common positive score: <most common positive score>, most common negative score: <most common negative score> LaMP-4 Given this author’s previous articles, try to describe a template for their headlines. I want to be able to accurately predict the headline gives one of their articles. Be specific about their style and wording, don’t tell me anything generic. Use the following format: The template is: ’[template 1]’, ’[template 2]’, ’[template 3]’, ’[template 4]’
 PAG Prompt Demo for LaMP-1
 Write a summary, in English, of the research interests and topics
 of a researcher who has published the following papers.
 Only generate the summary, no other text.
 The published papers are:
     \"Efficient Evaluation of Continuous Text Search Queries\",
     and \"Continuous Monitoring of Spatial Queries in Wireless
     Broadcast Environments\", and \"Spatial queries in wireless
     broadcast environments\", and \"Maximum Rank Query\", and
     \"Anonymous Query Processing in Road Networks\",
     and \"An Incremental Threshold Method for Continuous Text
     Search Queries\", and \"Continuous Top-k Monitoring on
     Document Streams.\", and \"Best upgrade plans for large road
     networks\", and \"Scalable verification for outsourced dynamic
     databases\", and \"Heuristic algorithms for balanced multi-way
     number partitioning\", and \"Aggregate nearest neighbor
     queries in spatial databases\", and \"Partially materialized
     digest scheme: an efficient verification method for outsourced
     databases\", and \"Best upgrade plans for single and multiple
     source-destination pairs.\", and \"Tree-based partition
     querying: a methodology for computing medoids in large spatial
     datasets\", and \"A Threshold-Based Algorithm for Continuous
     Monitoring of k Nearest Neighbors\", and \"Computing immutable
     regions for subspace top-k queries\", and \"Historical traffic-
     tolerant paths in road networks\"...
 PAG Prompt Demo for LaMP-2M
 Look at the following past movies this user has watched and determine
 the most popular tag they labeled. Answer in the following form:
 most popular tag: <tag top1>, <tag top2>, ..., <tag topn>.
 The movies and tags are:
     the tag for the movie: \"Young hobbit Frodo Baggins,
     after inheriting a mysterious ring from his uncle Bilbo,
     must leave his home in order to keep it from falling into
     the hands of its evil creator. Along the way,
     a fellowship is formed to protect the ringbearer
     and make sure that the ring arrives at its final destination:
     Mt. Doom, the only place where it can be destroyed.\" is
     \"fantasy\" , and the tag for the movie: \"Set in the 22nd
     century, The Matrix tells the story of a computer hacker
     who joins a group of underground insurgents fighting the vast
     and powerful computers who now rule the earth.\" is \"sci-fi\" ,
     and the tag for the movie: \"Batman raises the stakes in his
     war on crime. With the help of Lt. Jim Gordon and District
     Attorney Harvey Dent, Batman sets out to dismantle the
     remaining criminal organizations that plague the streets. The
     partnership proves to be effective, but they soon find
     themselves prey to a reign of chaos unleashed by a rising
     criminal mastermind known to the terrified citizens of Gotham
     as the Joker.\" is \"psychology\" , and the tag for the movie:
     \"An unsuspecting, disenchanted man finds himself working as a
     spy in the dangerous, high-stakes world of corporate
     espionage. Quickly getting way over-his-head, he teams up with
     a mysterious femme fatale.\" is \"twist ending\"...
 PAG Prompt Demo for LaMP-2N
 Look at the following past articles this journalist has
 written and determine the most popular category they write in.
 Answer in the following format: most popular category:
 <category top1>, <category top2>, ..., <category topn>.
 The articles and categories are:
     the category for the article:
     \"Champions like Tiger Woods are always charting and changing
     their course to be certain everything is on track. Tiger
     didn’t just come to Augusta because it was the popular thing
     to do. He wouldn’t have showed up if he wasn’t ready to win.
     He came to win and he’s prepared to win.\" is \"sports\" , and
     the category for the article: \"In 2011, in an interview with
     The Golf Channel, I predicted a Tiger Woods comeback while
     many others said he was done. I was right that time and I am
     right again, and I’ll say it right now and on the record:
     Tiger Woods will be back again and dominate the game of golf
     like the Tiger of old.\" is \"sports\" , and the category for
     the article: \"What do you teach your kids about money,
@@
     parents, the answer is probably\" is \"business\" , and the
     category for the article: \"With a little bit of planning and
     a lot of discipline, accomplishing your goals in the New Year
     can become a reality.  Imagine the immense satisfaction you’ll
     feel at this same time next year when you can look back and
     look at how far you’ve come and all that you have
     accomplished.\" is \"healthy living\" , and the category for
     the article: \"This whole argument boils down to a simple
     premise: who is in charge of our lives? Doctors? Politicians?
     Religious leaders? Or Us? Are we so feeble minded that we
     cannot be trusted to be responsible for our own existence?\"
     is \"politics\"...
 PAG Prompt Demo for LaMP-4
 Given this author’s previous articles, try to describe a
 template for their headlines. I want to be able to accurately
 predict the headline gives one of their articles. Be specific
 about their style and wording, don’t tell me anything generic.
 Use the following format: The template is: ’[template 1]’,
 ’[template 2]’, ’[template 3]’, ’[template 4]’.
 Previous articles and titles are:
     \"Selling a House to Buy a House\" is the title for \"Homeowners
     sell their homes and buy other homes for a variety
     of reasons including a need to live closer
     to a place of employment, to be closer to family, to enjoy a
     better climate, or simply to upgrade. This article is about
     finding the best sequence of steps in the process.\", and
     \"Investing In a Larger Down Payment: High Yields and No
     Risk\" is the title for \"Consumers looking to purchase a home
     within the near future face many decisions, including how
     large a down payment to make. The down payment is the sale
     price (confirmed by a appraisal) less the loan amount. In most
     cases, home purchasers must have financial assets at least as
     large as the down payment they make.\", and \"Why and How to
     Eliminate Mortgage Charges by Third Parties\" is the title for
     \"Third-party settlement costs could be eliminated by
     implementation of one simple rule: any service required by
     lenders as a condition for the granting of a home mortgage
     must be purchased and paid for by the lender.\", and \"Do Home
     Buyers Need a Pre-Approval?\" is the title for \"With
     bargaining power shifting from home buyers to sellers in an
     increasing number of local markets, buyers in competition with
     other buyers are looking for any edge they can get. One
     possible edge is a pre-approval letter (henceforth PAL) from a
     lender.\", and \"A New Challenge to the HECM Reverse Mortgage
     Program\" is the title for \"The United States today faces a
     retirement funds crisis: a rapidly growing number of persons
     who are retiring without the financial capacity to support
     themselves during ever-increasing life spans.\"...

### v2

@@ H.2 GSM-Hard
For M-Pilot, we employ a three-shot prompt to guide the decomposition steps, where <question> represents the current problem.
@@
Multi-turn Decomposition.
If the Black-box LLM encounters an error while solving the problem, M-Pilot is required to reflect on the previous decomposition <decompose> and re-decompose the problem.
@@
Code Generation.
Given the problem and the decomposition provided by M-Pilot, the Black-box model generates the corresponding code block for each sub-problem. We continue to use a three-shot prompt to instruct the Black-box model on how to translate the sub-problems into code, where <question> represents the current problem and <decompose> represents the decomposition provided by M-Pilot.
@@
<refinement> Prompt Lets use python to solve math problems. Here are three successful cases on how to do it, Q: Olivia has \$23. She bought five bagels for \$3 each. How much money does she have left? “‘ def solution(): ”””Olivia has \$23. She bought five bagels for \$3 each. How much money does she have left?””” money_initial = 23 bagels = 5 bagel_cost = 3 money_spent = bagels * bagel_cost money_left = money_initial - money_spent result = money_left return result “‘ Q: Michael had 58 golf balls. On tuesday, he lost 23 golf balls. On wednesday, he lost 2 more. How many golf balls did he have at the end of wednesday? “‘ def solution(): ”””Michael had 58 golf balls. On tuesday, he lost 23 golf balls. On wednesday, he lost 2 more. How many golf balls did he have at the end of wednesday?””” golf_balls_initial = 58 golf_balls_lost_tuesday = 23 golf_balls_lost_wednesday = 2 golf_balls_left = golf_balls_initial - golf_balls_lost_tuesday - golf_balls_lost_wednesday result = golf_balls_left return result “‘ Q: There were nine computers in the server room. Five more computers were installed each day, from monday to thursday. How many computers are now in the server room? “‘ def solution(): ”””There were nine computers in the server room. Five more computers were installed each day, from monday to thursday. How many computers are now in the server room?””” computers_initial = 9 computers_per_day = 5 num_days = 4 # 4 days between monday and thursday computers_added = computers_per_day * num_days computers_total = computers_initial + computers_added result = computers_total return result “‘ # Here is the actual question. Q: <question> You have generated code of solution() to solve the task. However, you executed the solution() function and get an error message: <error_msg> Referring to the successful case and the error message, you should complete the solution function with the correct code.
H.3 LaMP
Following the RAG-based framework [34] and the PAG-based framework [32], we implement prompt designs for both M-Pilot and the baseline methods. The prompt design for RAG is presented in Table 21, while the prompts for the two-stage PAG and M-Pilot are shown in Table 22. We create prompts for the controller model using the templates from Table 22 and subsequently combine the intermediate generations with the input question to form prompts for the environment model. Since LaMP-3 prompts are particularly lengthy, we provide additional examples of our PAG prompts for LaMP-1, LaMP-2N, LaMP-2M, and LaMP-4 as follows.
@@ H.3 LaMP
Task Per Profile Entry Prompt (PPEP) Aggregated Input Prompt (AIP) LaMP-1 ”Pi[title]” concat([PPEP(P1), …, PPEP(Pn)], ”, and ”). [INPUT] LaMP-2N ”the category for the article: ”Pi[text]” is ””Pi[category]”” concat([PPEP(P1), …, PPEP(Pn)], ”, and ”). [INPUT] LaMP-2M ”the tag for the movie: ”Pi[description]” is ”Pi[tag]” concat([PPEP(P1), …, PPEP(Pn)], ”, and ”). [INPUT] LaMP-3 Pi[score] is the score for ”Pi[text]” concat([PPEP(P1), …, PPEP(Pn)], ”, and ”). [INPUT] LaMP-4 ”Pi[title]” is the title for ”Pi[text]” concat([PPEP(P1), …, PPEP(Pn)], ”, and ”). [INPUT]
@@
Task Prompt LaMP-1 Write a summary, in English, of the research interests and topics of a researcher who has published the following papers. Only generate the summary, no other text. LaMP-2N Look at the following past articles this journalist has written and determine the most popular category they write in. Answer in the following format: most popular category: ¡category top1¿, ¡category top2¿, …, ¡category topn¿ LaMP-2M Look at the following past movies this user has watched and determine the mostpopular tag they labeled. Answer in the following form: most popular tag: ¡tag top1¿, ¡tag top2¿, …, ¡tag topn¿ LaMP-3 Based on this user’s past reviews, what are the most common scores they give for positive and negative reviews? Answer in the following form: most common positive score: ¡most common positive score¿, most common negative score: ¡most common negative score¿ LaMP-4 Given this author’s previous articles, try to describe a template for their headlines. I want to be able to accurately predict the headline gives one of their articles. Be specific about their style and wording, don’t tell me anything generic. Use the following format: The template is: ’[template 1]’, ’[template 2]’, ’[template 3]’, ’[template 4]’
PAG Prompt Demo for LaMP-1 Write a summary, in English, of the research interests and topics of a researcher who has published the following papers. Only generate the summary, no other text. The published papers are: \”Efficient Evaluation of Continuous Text Search Queries\”, and \”Continuous Monitoring of Spatial Queries in Wireless Broadcast Environments\”, and \”Spatial queries in wireless broadcast environments\”, and \”Maximum Rank Query\”, and \”Anonymous Query Processing in Road Networks\”, and \”An Incremental Threshold Method for Continuous Text Search Queries\”, and \”Continuous Top-k Monitoring on Document Streams.\”, and \”Best upgrade plans for large road networks\”, and \”Scalable verification for outsourced dynamic databases\”, and \”Heuristic algorithms for balanced multi-way number partitioning\”, and \”Aggregate nearest neighbor queries in spatial databases\”, and \”Partially materialized digest scheme: an efficient verification method for outsourced databases\”, and \”Best upgrade plans for single and multiple source-destination pairs.\”, and \”Tree-based partition querying: a methodology for computing medoids in large spatial datasets\”, and \”A Threshold-Based Algorithm for Continuous Monitoring of k Nearest Neighbors\”, and \”Computing immutable regions for subspace top-k queries\”, and \”Historical traffic- tolerant paths in road networks\”…
PAG Prompt Demo for LaMP-2M Look at the following past movies this user has watched and determine the most popular tag they labeled. Answer in the following form: most popular tag: <tag top1>, <tag top2>, …, <tag topn>. The movies and tags are: the tag for the movie: \”Young hobbit Frodo Baggins, after inheriting a mysterious ring from his uncle Bilbo, must leave his home in order to keep it from falling into the hands of its evil creator. Along the way, a fellowship is formed to protect the ringbearer and make sure that the ring arrives at its final destination: Mt. Doom, the only place where it can be destroyed.\” is \”fantasy\” , and the tag for the movie: \”Set in the 22nd century, The Matrix tells the story of a computer hacker who joins a group of underground insurgents fighting the vast and powerful computers who now rule the earth.\” is \”sci-fi\” , and the tag for the movie: \”Batman raises the stakes in his war on crime. With the help of Lt. Jim Gordon and District Attorney Harvey Dent, Batman sets out to dismantle the remaining criminal organizations that plague the streets. The partnership proves to be effective, but they soon find themselves prey to a reign of chaos unleashed by a rising criminal mastermind known to the terrified citizens of Gotham as the Joker.\” is \”psychology\” , and the tag for the movie: \”An unsuspecting, disenchanted man finds himself working as a spy in the dangerous, high-stakes world of corporate espionage. Quickly getting way over-his-head, he teams up with a mysterious femme fatale.\” is \”twist ending\”…
@@
PAG Prompt Demo for LaMP-4 Given this authors previous articles, try to describe a template for their headlines. I want to be able to accurately predict the headline gives one of their articles. Be specific about their style and wording, dont tell me anything generic. Use the following format: The template is: ’[template 1]’, ’[template 2]’, ’[template 3]’, ’[template 4]’. Previous articles and titles are: \”Selling a House to Buy a House\” is the title for \”Homeowners sell their homes and buy other homes for a variety of reasons including a need to live closer to a place of employment, to be closer to family, to enjoy a better climate, or simply to upgrade. This article is about finding the best sequence of steps in the process.\”, and \”Investing In a Larger Down Payment: High Yields and No Risk\” is the title for \”Consumers looking to purchase a home within the near future face many decisions, including how large a down payment to make. The down payment is the sale price (confirmed by a appraisal) less the loan amount. In most cases, home purchasers must have financial assets at least as large as the down payment they make.\”, and \”Why and How to Eliminate Mortgage Charges by Third Parties\” is the title for \”Third-party settlement costs could be eliminated by implementation of one simple rule: any service required by lenders as a condition for the granting of a home mortgage must be purchased and paid for by the lender.\”, and \”Do Home Buyers Need a Pre-Approval?\” is the title for \”With bargaining power shifting from home buyers to sellers in an increasing number of local markets, buyers in competition with other buyers are looking for any edge they can get. One possible edge is a pre-approval letter (henceforth PAL) from a lender.\”, and \”A New Challenge to the HECM Reverse Mortgage Program\” is the title for \”The United States today faces a retirement funds crisis: a rapidly growing number of persons who are retiring without the financial capacity to support themselves during ever-increasing life spans.\”…